# Hierarchical Sequence Iteration for Heterogeneous Question Answering

### v1

@@ Abstract
Retrieval-augmented generation (RAG) remains brittle on multi-step questions and heterogeneous evidence sources, trading accuracy against latency and token/tool budgets. This paper introduces Hierarchical Sequence (HSEQ) Iteration for Heterogeneous Question Answering, a unified framework that (i) linearize documents, tables, and knowledge graphs into a reversible hierarchical sequence with lightweight structural tags, and (ii) perform structure-aware iteration to collect just-enough evidence before answer synthesis. A Head Agent provides guidance that leads retrieval, while an Iteration Agent selects and expands HSeq via structure-respecting actions (e.g., parent/child hops, table row/column neighbors, KG relations); Finally the head agent composes canonicalized evidence to genearte the final answer, with an optional refinement loop to resolve detected contradictions. Experiments on HotpotQA (text), HybridQA/TAT-QA (table+text), and MetaQA (KG) show consistent EM/F1 gains over strong single-pass, multi-hop, and agentic RAG baselines with high efficiency. Besides, HSEQ exhibits three key advantages: (1) a format-agnostic unification that enables a single policy to operate across text, tables, and KGs without per-dataset specialization; (2) guided, budget-aware iteration that reduces unnecessary hops, tool calls, and tokens while preserving accuracy; and (3) evidence canonicalization for reliable QA, improving answers consistency and auditability.
@@ 1 Introduction
$\mathbf{C_{2}}:\textbf{Uncontrolled iteration and latency in multi-agent systems}$: With multi-agent collaboration and reasoning, agentic systems (Liu et al., 2025; Yang et al., 2025; Chen et al., 2025) easily explode the search space and can achieve multi-step reasoning. However they may fall with branchy plans, repeated web/file calls, and verbose chain-of-thought prompts, yielding unpredictable token/tool costs and latency; termination is often heuristic, leading to premature answers or extra wasted loops with budgets decoupled from the evidence actually inspected (Singh et al., 2025).
$\mathbf{C_{3}}:\textbf{Heterogeneity across formats}$: Free text, relational tables, and KGs typically require distinct indices, retrievers, prompt styles, and controller logic, preventing policy reuse and complicating training and deployment. Although existing heterogeneous RAG systems (Yu, 2022; Christmann & Weikum, 2024) are available to deal with multiple formats of data, they may still face issues in either weak alignment across representations or lossy and non-reversible serialization that obscures provenance and blocks faithful reconstruction.
Hierarchical Sequence Iteration (HSEQ) for Heterogeneous Question Answering introduces a reversible hierarchical sequence interface that linearizes documents, tables, and KGs into a sequence of typed segments with lightweight structure (e.g., parent/child locality, offsets or coordinates, minimal schema/time tags). An iteration policy operates on this unified substrate using short, budgeted steps: at each step it selects a few promising segments and predicts whether the accumulated set is sufficient to answer. A concise guidance plan—produced by a lightweight planner or a heuristic template—acts as a soft prior over which regions to probe first and when to stop. Once sufficiency is predicted, the selected segments are canonicalized into a compact, provenance-preserving package consumed by a head module to produce the final answer; an optional verifier can trigger a brief refinement if contradictions are detected.
To address above issues, this paper introduces HSEQ, a Hierarchical Sequence Iteration System that first recasts heterogeneous knowledge source into a single, LLM-native interface, then turning retrieval into a guided, budget-aware iterative process. The reversible HSEQ interface linearizes documents, tables, and KGs into a sequence of typed segments with lightweight structure (e.g., parent/child locality, offsets or coordinates, minimal schema/time tags). An iteration policy operates on this unified substrate using short, budgeted steps: at each step it selects a few promising segments and predicts whether the accumulated set is sufficient to answer. A concise guidance plan—produced by a lightweight planner or a heuristic template—acts as a soft prior over which regions to probe first and when to stop. Once sufficiency is predicted, the selected segments are canonicalized into a compact, provenance-preserving package consumed by a head module to produce the final answer; an optional verifier can trigger a brief refinement if contradictions are detected. pecifically, our key contributions are as followed:
  • Unified, reversible interface. A hierarchical sequence representation that standardizes text, tables, and KGs with lightweight structure and provenance, enabling a single controller to operate across formats.
@@
2 HSEQ: A Multi-Agent Heterogeneous Question Answering Framework
@@ 2.1 Background and Setup
Figure 1: HSEQ overview. (i) HSEQ-A linearizes heterogeneous sources into $S_{h}$ with level tags, parent pointers, and standardized metadata; (ii) HSEQ-I iterates over a windowed stream of segments under budgets, guided by $g$, and queries $\Phi$ for sufficiency; (iii) $\kappa$ compacts $M_{t}$ into provenance-preserving evidence; (iv) HSEQ-H produces the final answer and optionally triggers a brief refinement if inconsistencies are detected.
2.2 HSEQ Architecture
The proposed system couples a unified hierarchical sequence(HSEQ) representation with an iteration policy and a head module $\mathcal{H}$ for answer synthesis. Let $q$ denote a user query and $D$ a heterogeneous corpus. A modality-aware adapter $\tau$ converts $D$ into a single hierarchical sequence $S_{h}=\tau(D)$. An iteration module $\pi_{\theta}$ operates on $(q,S_{h})$ and maintains an evolving evidence set $M_{t}$. The final evidence $M^{\star}$ is then canonicalized by $\kappa$ and passed to $\mathcal{H}$ for final answer generation. The end-to-end mapping is summarized as
$F=\big(\tau,\pi_{\theta},\Phi,\kappa,\mathcal{H}\big),\qquad F(q,D)\;=\;\mathcal{H}\!\left(q,\,\kappa\!\left(M^{\star}\right)\right),$ (1)
Specifically, during iteration module $\pi_{\theta}$ selecting and expanding segments on $S_{h}$, the budget-aware sufficiency criterion $\Phi$ and the budget state $B_{t}$ (tokens, tool calls, steps) functioned inside the module to decide when the accumulated evidence is adequate for answering as well as triggering potential early stopping.
@@ 2.2 HSEQ Architecture
Generally, to achieve iteration through an unified data structure building from heterogeneous data sources, the HSEQ framework consists of three key modules: HSEQ-Adapter (HSEQ-A), HSEQ-Iterator (HSEQ-I), and HSEQ-Head (HSEQ-H).
@@ 2.3 HSEQ-Adapter(HSEQ-A)
The HSEQ-Adapter is build to produce unified structure(HSEQ $S_{h}$) that exposes locality (parent/child), alignment (span or coordinates), and lightweight semantics (time, schema, language) in a modality-agnostic format, while remaining reconstructable. Formally, each item $x_{j}$ is mapped by a modality-specific adapter $\tau_{m_{j}}$ to a finite set of segments $\tau_{m_{j}}(x_{j})\subset\mathcal{S}$ and then concatenated:
$S_{h}\;=\;\bigsqcup_{j=1}^{N}\tau_{m_{j}}(x_{j})\;\in\;\mathcal{S}^{\ast},\quad\mathcal{S}\ni s\;=\;\big(\mathrm{id}(s),\,\ell(s),\,p(s),\,c(s),\,\mu(s)\big).$ (3)
Figure 2: HSEQ $S_{h}$ construction: Different modalities of data are transformed into unified sequence by HSEQ-A
Here $\ell(s)$ is a level tag matching the raw content, including sentence, paragraph, table, triplet, etc., while $p(s)$ is a parent pointer recording the roots. $c(s)$ is compact human-readable content, and $\mu(s)$ is metadata with fixed keys to record content attributes.
The single, modality-aware adapter converts heterogeneous sources into a common sequence of hierarchical segments. After the construction, each segment is a lightweight record $s=\big(\textit{id,\,level,\,parent,\,content,\,metadata}\big)$, where level marks granularity (e.g., document/paragraph/sentence, table_row/table_cell, triplet/subgraph), parent keeps locality via a unique container pointer, content is a compact human-readable summary (text span, serialized row, or compact triple), and metadata standardizes provenance and alignment with fixed keys (e.g., source_id, uri, offsets, schema, time). Segments are concatenated into a final usable $S_{h}$ in parent-before-child order. This minimal contract enables structure-aware neighborhoods and budget-aware iteration without inspecting raw files.
2.4 HSEQ-Iterator(HSEQ-I)
After HSEQ $S_{h}$ are build, the HSEQ-Iterator $\pi_{\theta}$ can be used on $(q,S_{h})$ and maintains an evolving evidence set $M_{t}$ regarding question $q$.
@@ 2.4 HSEQ-Iterator(HSEQ-I)
Let $M_{t}\subseteq S_{h}$ denote the selected evidence at step $t$, $C_{t}\subseteq S_{h}$ a candidate window obeying a budget state $B_{t}$, and $\mathcal{N}(\cdot)$ the structure-aware neighborhood operators induced by levels, parents, and coordinates. The HSEQ-I module $\pi_{\theta}$ functions each step following the policy
@@
2.5 HSEQ-Head (HSEQ-H).
The HSEQ-Head module $\mathcal{H}$ can be used in two parts: 1) Guiding the retrieval for HSEQ-I; and 2) Generating final conclusion regarding the question.
@@
Algorithm 1 Guided Iterative Selection under HSEQ-I
@@
3 Learning to Use HSEQ with Open-Source LLMs
3.1 Fine-tuning HSEQ-I
Although HSEQ-I can be directly called using unfinetuned LLMs, a supervised finetuning is needed to facilitate the iteration and improve LLM’s understandability with HSEQ.
Figure 3: HSEQ_I is build by training questions from multiple datasets. After guidance sets are generated, LoRA is applied for finetuning
@@ 3.1 Fine-tuning HSEQ-I
Supervision is organized as tuples $(q,\;\mathrm{type},\;S_{h},\;A^{\star})$. Beside above mentioned query $q$ and HSEQ $S_{h}$, an optional question label $\mathrm{type}$ is added during training. To ensure Iteration module $\pi_{\theta}$ can trace sufficient data while limiting usage, a target trajectory $A^{\star}=\{(a_{t}^{\star},s_{t}^{\star})\}_{t=1}^{\tau^{\star}}$ is added which consists of an action $a_{t}^{\star}$ and a binary sufficiency signal $s_{t}^{\star}\in\{0,1\}$, with stopping time $\tau^{\star}=\min\{t:\,s_{t}^{\star}=1\}$. When explicit trajectories are unavailable, weak positives $P^{\star}\subseteq S_{h}$ are induced by high-precision matching between the gold answer (or oracle spans) and segment content, optionally augmented by lexical overlap with $q$. A target action sequence is then synthesized by greedily selecting from $P^{\star}$ under the budget (details in App. A.2).
@@
Let the step state be $(q,S_{h},M_{t},C_{t},g,B_{t})$. The policy is trained by supervised risk minimization over trajectories with a parameter-efficient adaptation of a base LLM. With teacher forcing for $t<\tau^{\star}$,
$\min_{\theta}\ \mathbb{E}\bigg[\sum_{t=1}^{\tau^{\star}}\underbrace{\ell_{\mathrm{act}}\!\left(\pi_{\theta}(\cdot\mid\mathrm{state}_{t}),\,a_{t}^{\star}\right)}_{\text{action loss}}\;+\;\lambda\,\underbrace{\ell_{\mathrm{stop}}\!\left(\pi_{\theta}(\cdot\mid\mathrm{state}_{t}),\,s_{t}^{\star}\right)}_{\text{sufficiency loss}}\bigg],$
where $\mathrm{state}_{t}=(q,S_{h},M_{t},C_{t},g,B_{t})$ and $\lambda>0$ balances early stopping. When $A^{\star}$ is synthesized from $P^{\star}$, per-step weights attenuate low-confidence choices to reduce label noise (App. A.2). During experiments, low-Rank Adaptaion is used for model finetuning (Hu et al., 2022) (App. A.3.4)
3.2 HSEQ-H: Guidance and Answer Generation
Guidance generation (HSEQ-H).
Given a question $q$ (and optional type tag), HSEQ-H produces a short guidance $g$ that steers the iterator toward promising parts of $S_{h}$ and specifies a stop condition. We use two interchangeable modes: (i) a lightweight planner that drafts $g$ in 2–4 sentences, and (ii) a heuristic template keyed by coarse question patterns (e.g., number/factoid/yes–no). Each $g$ follows the same simple structure: (first-look targets) what to check first (e.g., entities/rows/1–2-hop neighbors); (expansion rule) how to broaden if evidence is missing (parent/child, row/column, or relation hops); (stop rule) when to stop (e.g., “answer span/number is explicit and corroborated by $\geq 1$ snippet”). Guidance is cached per example and reused at inference; on a cache miss, HSEQ-H generates $g$ (planner) or falls back to the template. Importantly, $g$ is a soft prior: the iterator may override it when stronger signals appear.
Answer generation (HSEQ-H).
After the iterator halts with selected evidence $M_{\tau}$, a canonicalizer $\kappa$ compacts it into a short, provenance-preserving package $Z=\kappa(M_{\tau})$ (snippet text plus ids/levels/source and minimal offsets/coordinates). HSEQ-H then performs evidence-conditioned answering:
$\hat{y}\;=\;\mathcal{H}\!\big(q,\,Z;\,g\big),$
where the prompt is kept minimal: answer only (span/number/yes–no), grounded in $Z$, no chain-of-thought. When useful, HSEQ-H also returns supporting ids from $Z$ for auditability. An optional lightweight check (e.g., “does some snippet in $Z$ entail $\hat{y}$?”) can trigger a one-shot refinement—the iterator resumes for a few steps under a tightened $g^{\prime}$—otherwise $\hat{y}$ is emitted. This setting aligns answers with explicit evidence, and preserves the modality-agnostic contract of HSEQ.

### v3

# RELOOP: Recursive Retrieval with Multi-Hop Reasoner and Planners for Heterogeneous QA

@@ Abstract
Retrieval-augmented generation (RAG) remains brittle on multi-step questions and heterogeneous evidence sources, trading accuracy against latency and token/tool budgets. This paper introduces RELOOP, a structure aware framework using Hierarchical Sequence (HSEQ) that (i) linearize documents, tables, and knowledge graphs into a reversible hierarchical sequence with lightweight structural tags, and (ii) perform structure-aware iteration to collect just-enough evidence before answer synthesis. A Head Agent provides guidance that leads retrieval, while an Iteration Agent selects and expands HSeq via structure-respecting actions (e.g., parent/child hops, table row/column neighbors, KG relations); Finally the head agent composes canonicalized evidence to genearte the final answer, with an optional refinement loop to resolve detected contradictions. Experiments on HotpotQA (text), HybridQA/TAT-QA (table+text), and MetaQA (KG) show consistent EM/F1 gains over strong single-pass, multi-hop, and agentic RAG baselines with high efficiency. Besides, RELOOP exhibits three key advantages: (1) a format-agnostic unification that enables a single policy to operate across text, tables, and KGs without per-dataset specialization; (2) guided, budget-aware iteration that reduces unnecessary hops, tool calls, and tokens while preserving accuracy; and (3) evidence canonicalization for reliable QA, improving answers consistency and auditability. Codes are available at https://anonymous.4open.science/r/RELOOP
RELOOP: Recursive Retrieval with Multi-Hop Reasoner and Planners for Heterogeneous QA
Ruiyi Yang1, Hao Xue1,2, Imran Razzak1,3, Hakim Hacid4, Flora D. Salim1, 1University of New South Wales, 2The Hong Kong University of Science and Technology (Guangzhou), 3Mohamed Bin Zayed University of Artificial Intelligence, 4Technology Innovation Institute, Correspondence: ruiyi.yang@unsw.edu.au
@@ 1 Introduction
$\mathbf{C_{2}}:\textbf{Uncontrolled iteration and latency}$: With multi-agent collaboration and reasoning, agentic systems (Liu et al., 2025; Yang et al., 2025; Chen et al., 2025) easily explode the search space and can achieve multi-step reasoning. However they may fall with branchy plans, repeated web/file calls, and verbose chain-of-thought prompts, yielding unpredictable token/tool costs and latency; termination is often heuristic, leading to premature answers or extra wasted loops with budgets decoupled from the evidence actually inspected (Singh et al., 2025).
$\mathbf{C_{3}}:\textbf{Heterogeneity across formats}$: Free text, relational tables, and KGs typically require distinct indices, retrievers, prompt styles, and controller logic, preventing policy reuse and complicating training and deployment. Although existing heterogeneous RAG systems (Yu, 2022; Christmann and Weikum, 2024) are available to deal with multiple formats of data, they may still face issues in either weak alignment across representations or lossy and non-reversible serialization that obscures provenance and blocks faithful reconstruction.
To address above issues, this paper introduces RELOOP, a Hierarchical Sequence Iteration System that first recasts heterogeneous knowledge source into a single, LLM-native interface, then turning retrieval into a guided, budget-aware iterative process. The reversible structured parsing linearizes documents, tables, and KGs into a sequence of typed segments with lightweight structure (e.g., parent/child locality, offsets or coordinates, minimal schema/time tags). An iteration policy operates on this unified substrate using short, budgeted steps: at each step it selects a few promising segments and predicts whether the accumulated set is sufficient to answer. A concise guidance plan—produced by a lightweight planner or a heuristic template—acts as a soft prior over which regions to probe first and when to stop. Once sufficiency is predicted, the selected segments are canonicalized into a compact, provenance-preserving package consumed by a head module to produce the final answer; an optional verifier can trigger a brief refinement if contradictions are detected. pecifically, our key contributions are as followed:
  • Unified, reversible structured parsing. A hierarchical sequence representation(HSEQ) that standardizes text, tables, and KGs with lightweight structure and source, enabling a single parser to operate across formats.
@@
2 Related Work
LLM Finetuning
Large Language Models (LLMs) often adopt finetuning to unlock their capabilities for downstream applications, like medical (Goyal et al., 2024), economic Guo and Yang (2024), or human activity recognition Li et al. (2024). To enhance finetuning efficiency, methods like quantization (Dettmers et al., 2022) parameter efficient fine tuning (Hu et al., 2022; Dettmers et al., 2023; Li and Liang, 2021) can be applied.
Retrieval Augmented Generation
RAG systems help LLMs retrieve extra knowledge according to queries and thereby improving the accuracy of LLM response (Fan et al., 2024), with no necessity to finetune the model. External databases ensure knowledge offered is domain-specific and timely, adding reliability and interpretability (Lewis et al., 2020; Jiang et al., 2023). Accuracy of knowledge retrieval and quality of responses are two key factors for RAG systems evaluation (Yu et al., 2024). Apart from text, table, or html sources (Guo et al., 2024b; Chan et al., 2024; Jin et al., 2025), recent researches have combined graph-structured data into RAG systems(GraphRAG) to improve the efficiency of knowledge interpretability by capturing relationships between entities and utilizing triplets as the primary data source (Edge et al., 2024; Peng et al., 2024; Hu et al., 2024; Mavromatis and Karypis, 2024).
Multi Agent QA system
Structural and unified RAG interfaces.
Beyond standard text-centric RAG, a line of work introduces structural or unified retrieval layers. Graph-based RAG systems construct heterogeneous or chunk-level graphs where nodes represent passages, entities, or sections, and edges encode semantic or hyperlink connectivity; retrieval then propagates over this graph to improve multi-hop reasoning and global coverage (Wu et al., 2024; Huang et al., 2025; Luo et al., 2025). Other systems build hierarchical or modular indices over mixed document formats, or define unified data schemas for training language agents and tools, but still operate over opaque contexts at inference time (Reynolds and Corrigan, 2024; Liu et al., 2025; Chen et al., 2024).
While existing RAG-based methods still suffered from limitation mentioned above, there is a rising need for RAG interfaces that (i) preserve modality-specific structure in a reversible way rather than collapsing all sources into an opaque graph or index; (ii) expose a generic, LLM-native segment schema with explicit level, parent, and alignment fields so that a single controller can navigate text, tables, and KGs uniformly; and (iii) couple this interface with an explicit sufficiency-aware, budget-controlled selection process, so that evidence gathering is both auditable and aligned with resource constraints. RELOOP is designed to meet these requirements by treating all sources as typed, provenance-aware segments in a hierarchical sequence and pairing this representation with a learned sufficiency head and budget-aware iteration policy.
3 RELOOP: A Multi-Agent Heterogeneous Question Answering Framework
Figure 1: RELOOP overview. (i) HSEQ-A linearizes heterogeneous sources into HSEQ $S_{h}$ with level tags, parent pointers, and standardized metadata; (ii) RELOOP-I iterates over a windowed stream of segments under budgets, guided by $g$, and queries $\Phi$ for sufficiency; (iii) $\kappa$ compacts $M_{t}$ into provenance-preserving evidence; (iv) RELOOP-H produces the final answer and optionally triggers a brief refinement if inconsistencies are detected.
@@ 3.1 Background and Setup
3.2 RELOOP Architecture
The proposed system couples a unified hierarchical sequence (HSEQ) representation with an iteration policy and a head module $\mathcal{H}$ for answer synthesis. As described in Section 3.1, a modality-aware adapter $\tau$ converts $D$ into a single hierarchical sequence $S_{h}$, and the iterator $\pi_{\theta}$ accumulates evidence under budgets $B$, with framework as:
$F=\big(\tau,\pi_{\theta},\Phi,\kappa,\mathcal{H}\big),\quad F(q,D)\;=\;\mathcal{H}\!\left(q,\,\kappa\!\left(M^{\star}\right)\right),$ (1)
where $\Phi$ denotes a budget-aware sufficiency criterion used by $\pi_{\theta}$, and $\kappa$ canonicalizes the final evidence $M^{\star}$ for $\mathcal{H}$.
During iteration, $\pi_{\theta}$ selects and expands segments on $S_{h}$ under the budget state $B_{t}$ and stops when evidence is deemed sufficient by $\Phi$ or the budget is exhausted:
@@ 3.2 RELOOP Architecture
Generally, to achieve iteration through an unified data structure building from heterogeneous data sources, the RELOOP framework consists of three key modules: HSEQ-Adapter (HSEQ-A), RELOOP-Iterator (RELOOP-I), and RELOOP-Head (RELOOP-H).
@@ 3.3 HSEQ-Adapter(HSEQ-A)
The HSEQ-Adapter is build to produce unified structure(HSEQ $S_{h}$) that exposes locality (parent/child), alignment (span or coordinates), and lightweight semantics (time, schema, language) in a modality-agnostic format, while remaining reconstructable. Formally, each item $x_{j}$ is mapped by a modality-specific adapter $\tau_{m_{j}}$ to a finite set of segments $\tau_{m_{j}}(x_{j})\subset\mathcal{S}$ and then concatenated as $S_{h}=\bigsqcup_{j=1}^{N}\tau_{m_{j}}(x_{j})\in\mathcal{S}^{\ast}$, where each $s\in\mathcal{S}$ is represented as $s=(\mathrm{id}(s),\ell(s),p(s),c(s),\mu(s))$.
$\ell(s)$ is a level tag matching the raw content, including sentence, paragraph, table, triplet, etc., while $p(s)$ is a parent pointer recording the roots. $c(s)$ is human-readable content, and $\mu(s)$ is metadata with fixed keys to record content attributes.
Segments are concatenated into a final usable $S_{h}$ in parent-before-child order. This minimal contract enables structure-aware neighborhoods and budget-aware iteration without inspecting raw files.
3.4 RELOOP-Iterator(RELOOP-I)
Following details how the iterator $\pi_{\theta}$ is guided by $g$ and controlled by windowing, neighborhood expansion, and a sufficiency-aware stopping signal.
@@ 3.4 RELOOP-Iterator(RELOOP-I)
Let $M_{t}\subseteq S_{h}$ denote the selected evidence at step $t$, $C_{t}\subseteq S_{h}$ a candidate window obeying a budget state $B_{t}$, and $\mathcal{N}(\cdot)$ the structure-aware neighborhood operators induced by levels, parents, and coordinates. The RELOOP-I module $\pi_{\theta}$ functions each step following the policy
@@
One-step worked example.
Suppose $\rho$ is paragraph $\prec$ row $\prec$ sentence $\prec$ triplet and W=5W\!=\!5. At t=0t\!=\!0, $C_{0}$ is the first $5$ segments under $\rho$. The policy selects $K_{1}\!\subseteq\!C_{0}$ ($|K_{1}|\!\leq\!k$) and optionally expands with $\mathcal{N}_{\text{children}}$ (to get sentences within a paragraph) or $\mathcal{N}_{\text{row}}$ (to fetch a full table row when a cell is promising). Then $M_{1}\!=\!M_{0}\cup K_{1}$. $\mathrm{Refresh}$ advances the window to the next $5$ unseen segments in $S_{h}$. If $\Phi$ deems evidence sufficient ($s_{1}\!=\!1$) and $t\!\geq\!T_{\min}$, iteration halts; otherwise proceed to t=2t\!=\!2 with updated $B_{t}$.
3.5 RELOOP-Head (RELOOP-H).
The RELOOP-Head module $\mathcal{H}$ can be used in two parts: 1) Guiding the retrieval for RELOOP-I; and 2) Generating final conclusion regarding the question.
@@
Algorithm 1 Guided Iterative Selection under RELOOP-I
@@
4 RELOOP usage with Open-Source LLMs
With Section 3 as the theoretical interface, this section details how to instantiate RELOOP with open-source LLMs.
4.1 Fine-tuning RELOOP-I
@@
We organize supervision as tuples $(q,\mathrm{type},S_{h},A^{\star})$. Besides $q$ and $S_{h}$, an optional label $\mathrm{type}$ is added. The trajectory $A^{\star}=\{(a_{t}^{\star},s_{t}^{\star})\}_{t=1}^{\tau^{\star}}$ contains an action and a binary sufficiency signal with $\tau^{\star}=\min\{t:\,s_{t}^{\star}=1\}$. When explicit trajectories are unavailable, weak positives $P^{\star}\subseteq S_{h}$ are induced by high-precision matching between gold answers (or oracle spans) and segment content, optionally augmented by lexical overlap with $q$. A target action sequence is synthesized by greedily selecting from $P^{\star}$ under the budget (details in App. A.2).
@@
The policy $\pi_{\theta}$ is trained with parameter-efficient adaptation of a base LLM using Low-Rank Adaptation (LoRA) (Hu et al., 2022) (App. A.3.4). With teacher forcing for $t<\tau^{\star}$, the objective minimizes an action-matching loss $\ell_{\mathrm{act}}$ and a sufficiency-aware stopping loss $\ell_{\mathrm{stop}}$, as formalized in Eq. 3, where $\mathrm{state}_{t}=(q,S_{h},M_{t},C_{t},g,B_{t})$ and $\lambda>0$ trades off action accuracy and early stopping. When $A^{\star}$ is synthesized from $P^{\star}$, per-step weights attenuate low-confidence choices to reduce label noise (App. A.2).
$\min_{\theta}\quad\mathbb{E}\Bigg[\sum_{t=1}^{\tau^{\star}}$ $\;\ell_{\mathrm{act}}\!\left(\pi_{\theta}(\cdot\mid\mathrm{state}_{t}),\,a_{t}^{\star}\right)$ (3)
$\;+\;\lambda\,\ell_{\mathrm{stop}}\!\left(\pi_{\theta}(\cdot\mid\mathrm{state}_{t}),\,s_{t}^{\star}\right)\Bigg].$
Figure 2: RELOOP-I is trained from multi-source questions. After guidance sets are prepared, LoRA is applied for finetuning.
Guidance generation (RELOOP-H).
Given a question $q$ (and optional $\mathrm{type}$), RELOOP-H produces a short guidance $g$ that steers the iterator and specifies a stop rule. We use two modes: (i) a lightweight planner that drafts $g$ in 2–4 sentences; (ii) a heuristic template keyed by coarse question patterns (e.g., number/factoid/yes–no). $g$ follows a fixed structure: first-look targets (entities/rows/1–2-hop neighbors), expansion rule (parent/child, row/column, or relation hops), and stop rule (e.g., “answer span/number is explicit and corroborated by $\geq 1$ snippet”). $g$ is cached and reused; on a cache miss, we generate or fall back to the template. $g$ is a soft prior—the iterator may override it when stronger signals appear.
Evidence-conditioned answering (RELOOP-H).
[…]